I need help..i don't understand to multiply and simplify
b^15 * b^0 =...wouldn't that be b^)

Answer by sudhanshu_kmr(1152)


b^15 * b^0
= b^(15 +0)

= b^15
and, also note that b^0 = 1
another way,
b^15 * 1
=b^15